# Multi-modal Misinformation Detection: Approaches, Challenges and Opportunities

### v6

@@ 3. Multi-modal model study
Extracted features and the way they are fused play an important role in the model architecture. In fact, the model-based and feature-based studies are closely related through fusion strategies which makes the demarcation of these two studies very difficult. Hence, in this section, first we discuss common fusion strategies as the point of connection between the two studies. Furthermore, we categorize existing works based on the machine learning techniques that are exploited by each work. Specifically, we classify them into two main categories: 1) classic machine learning, and 2) deep learning-based solutions. In this section, we discuss each category in detail.
(a) Early fusion mechanism.
(b) Late fusion mechanism.
(c) A hybrid of early and late fusion mechanisms.
Figure 2. Different fusion mechanisms in multi-modal learning.
@@ 3.1. Fusion mechanisms
Data fusion is the process of combining information from multiple modalities in order to take advantage of all different aspects of the data and extract as much information as possible to improve the performance of machine learning models as opposed to using a single data aspect or modality. Different fusion mechanisms have been used to combine features from different modalities, including those we mentioned in the previous section. Fusion mechanisms are often categorized into one of the following groups:
@@
also known as feature-level fusion refers to combining features from different data modalities at an early stage using an operation which is often concatenation. This type of fusion is often performed ahead of classification. If the fusion process is done after feature extraction, it is sometimes referred to as intermediate fusion (Martínez and Yannakakis, 2014; Lahat et al., 2015; Boulahia et al., 2021).
@@
also known as decision-level or kernel-level fusion, is usually done in the classification stage. This method thus depends on the results obtained by each data modality individually. In other words, the modality-wise classification results are combined using techniques such as sum, max, average, and weighted average. Most of the late fusion solutions use handcrafted rules, prone to human bias, and are far from real-world peculiarities (Martínez and Yannakakis, 2014; Lahat et al., 2015; Boulahia et al., 2021).
Comparison of fusion mechanisms
In most cases, early fusion is a complex operation, whereas late fusion is easier to perform (Atrey et al., 2010) because unlike the early fusion where the features from different modalities e.g., image and text may have different representation, the decisions at the semantic level usually have the same representation. Therefore, the fusion of decisions is easier than the fusion of features. However, the late fusion strategy does not utilize the feature level correlation among modalities which may improve the classification performance. In fact, it is shown that in many cases, the early fusion of different modalities, outperforms multi-modal late fusion while applying deep learning or classic machine learning classifiers (Dong et al., 2014; Gallo et al., 2020). For instance, early fusion of images and texts while using BERT and CNN, on UPMC Food-101 dataset [^8] http://visiir.lip6.fr/explore (Wang et al., 2015) outperforms late fusion of these modalities.
Another advantage of early fusion is that it requires less computation time because training is performed only once but the late fusion needs multiple classifiers for local decisions (Atrey et al., 2010). However, to have the best of both worlds, there are hybrid approaches as well which take advantage of both early and late fusion strategies (Atrey et al., 2010). A simplified scheme of different fusion mechanisms for multi-modal learning is demonstrated in Fig. 2. Traditional and modern approaches for detecting multi-modal misinformation detection, some of which employ fusion mechanisms, are covered in the section that follows.
@@ 3.2. Classic machine learning solutions
As we discussed earlier, a vast majority of misinformation detection leverage a single modality a.k.a. aspect of news articles e.g., text (Horne and Adali, 2017; Wu et al., 2017; Guacho et al., 2018; Shu et al., 2020a), image (Huh et al., 2018; Qi et al., 2019; Abdali et al., 2021a; Choudhary and Arora, 2021), user features (Wu and Liu, 2018; Shu et al., 2018, 2019c), and temporal properties (Kumar and Shah, 2018; Shu et al., 2019b; Song et al., 2021b). However, recently, there have been very few works that incorporate various aspects of a news article using classic machine learning techniques in order to create multi-modal article representations.
For instance, a work by Shu et al. (K. Shu and Liu, 2019), proposes individual embedded representations for text, user-user interactions, user-article interactions and publisher-article interactions, and defines a joint optimization problem leveraging these individual representations. Finally, they apply “Non-convex Optimization” solution via the Alternating Least Squares (ALS) algorithm to solve the proposed optimization problem.
@@
More recently, Meel et al. (Hakak et al., 2021), have proposed an “Ensemble Framework” which leverages text embedding, a score calculated by cosine similarity between image caption and news body, and noisy images. Despite the fact that some of the modules of this model e.g., text embedding generator leverage deep attention-based architecture, the classification process is done via a classic ensemble technique i.e. max “Voting”.
Summarily, due to the success of deep learning-based techniques in feature extraction and classification tasks, classic machine learning-based techniques are not commonly used these days. However, considering the fact that deep learning techniques are data hungry and require a lot of effort for training and fine-tuning the models, depending on the applications, they are still being used solely or in conjunction with deep learning techniques.
@@ 3.3. Deep learning solutions
Due to the impressive success of deep neural networks in feature extraction and classification of text, image and many other modalities, over the past few years, they have been widely exploited by research scientists for a variety of multi-modal tasks, including misinformation detection. We may categorize deep learning-based multi-modal misinformation detection into five categories: concatenation-based, attention-based, generative-based, graph neural network-based and cross-modality discordance-aware architectures. In what follows, we summarize and categorize the existing works into the aforementioned categories.
@@ 3.3.1. Concatenation-based architectures
The majority of the existing work on multi-modal misinformation detection, embed each modality e.g., text or image into a vector representation and then concatenate them to generate a multi-modal representation which can be utilized for classification tasks. For instance, Singhal et al. propose to use a pretrained XLnet and VGG-19 models to embed text and image respectively and then classify the concatenation of resulted feature vectors to detect misinformation (Singhal et al., 2019).
In another work (Alonso-Bartolome and Segura-Bedmar, 2021), Bartolome et al. exploit a Convolutional Neural Network (CNN) that takes as inputs both text and image corresponding to an article and the outputs are concatenated into a single vector. Qi et al. extract text, Optical Character Recognition (OCR) content, news-related high-level semantics of images e.g., celebrities and landmarks, and visual CNN features of the image and then in the stage of multi-modal features fusion, text-image correlations, mutual enhancement, and entity inconsistency are merged by concatenation operation (Qi et al., 2021).
In another work (Rezayi et al., 2021), Rezayi et al. leverage network, textual and relaying features such as hashtags and URLs and classify articles using the concatenation of the feature embeddings.Works in (Shu et al., 2019a; Raj and Meel, 2022) are other examples of this category of deep learning-based solutions.
@@ 3.3.2. Attention-based architectures
As mentioned above, many architectures simply concatenate vector representations, thereby fail to build effective multi-modal embeddings. Such models are not efficient in many cases. For instance, the entire text of an article does not necessarily need to be false for the corresponding image and vice versa to consider the article as misinformative content. Thus, some recent works attempt to use the attention mechanism to attend to relevant parts of image, text etc. Attention mechanism is a more effective approach for utilizing embeddings, as it produces richer multi-modal representations.
@@
For instance, a work by Sachan et al. (Sachan et al., 2021), proposes Shared Cross Attention Transformer Encoders (SCADE) which exploits CNNs and transformer-based methods to encode image and text information and utilizes cross-modal attention and shared layers for the two modalities. SCADE pays attention to the relevant parts of each modality with reference to the other one.
@@
Another example is a work by Kumari et al. (Kumari and Ekbal, 2021), where a framework is developed to maximize the correlation between textual and visual information. This framework has four different sub-modules. Attention Based Stacked Bidirectional Long Short Term Memory (ABS-BiLSTM) for textual feature representation, Attention Based Multilevel Convolutional Neural Network–Recurrent Neural Network (ABM-CNN–RNN) for visual feature extraction, multi-modal Factorized Bilinear Pooling (MFB) for feature fusion and finally Multi-Layer Perceptron (MLP) for the classification.
@@
In another work, Qian et al. (Qian et al., 2021), proposes a Hierarchical Multi-modal Contextual Attention Network (HMCAN) architecture which utilizes a pretrained Bidirectional Encoder Representations from Transformers (BERT) and convolutional ResNet50 to generate word and image embeddings and a multi-modal contextual attention network to explore the multi-modal context information. In this work, HMCAN leverages different multi-modal contextual attention networks to constitute a hierarchical encoding network to explore and capture the rich hierarchical semantics of multi-modal data.
@@
Another example is (Jin et al., 2017), where Jin et al. fuse features from three modalities i.e., textual, visual and social context using an RNN that utilizes attention mechanism (att-RNN) for feature alignment. Jing et al. propose TRANSFAKE (Jing et al., 2021) to connect features of text and images into a series, and feed them into a vision-language transformer model to learn the joint representation of multi-modal features. TRANSFAKE adopts a preprocessing method similar to BERT for concatenated text, comments and images.
@@
In another work (Wang et al., 2022), Wang et al. apply scaled dot-product attention on top of image and text features as a fine-grained fusion and use the fused feature to classify articles.
@@
Wang et al. propose a deep learning network for Biomedical informatics that leverages visual and textual information and a semantic- and task-level attention mechanism to focus on the essential contents of a post that signal anti-vaccine messages (Wang et al., 2021c).
@@
Another example is a work by Lu et al. where representations of user interaction, word representations and propagation features are concatenated after applying a dual co-attention mechanism. The rationale is to capture the correlations between users’ interaction/propagation and tweet’s text (Lu and Li, 2020).
@@
Finally, Song et al. (Song et al., 2021a), propose a multi-modal fake news detection architecture based on Cross-modal Attention Residual (CARN) and Multichannel convolutional neural Networks (CARMN). Cross-modal Attention Residual or CARN selectively extracts the information related to a target modality from a source modality while maintaining the unique information of the target.
@@ 3.3.3. Generative architectures
In this category of deep learning solutions, the goal is to either apply Generative Networks or use auxiliary networks to learn individual or multi-modal representations, spaces or parameters in order to improve the classification performance of the fake news detector.
@@
As an example, Jaiswal et al. propose a BERT based multi-modal variational Autoencoder (VAE) (Jaiswal et al., 2021) that consists of an encoder, decoder and a fake news detector. The encoder encodes the shared representations of both the image and text into a multidimensional latent vector. The decoder decodes the multidimensional latent vector into the original image and text and the fake news detector is a binary classifier that takes the shared representation as an input and classifies it as either fake or real.
@@
Similarly, Kattar et al. propose a deep multi-modal variational autoencoder (MVAE) (Khattar et al., 2019) which learns a unified representation of both the modalities of a tweet’s content. Similar to the previous work, MVAE has three main components: an encoder, a decoder and a fake News detector that utilizes the learned shared representation to predict if a news is fake or not.
@@
Like previous work, a work by Zeng et al. (Zeng et al., 2020) proposes to capture the correlations between text and image by a VAE-based multi-modal feature fusion method. In another work, Wang et al. propose Event Adversarial Neural Networks (EANN) (Wang et al., 2018), an end-to-end framework which can derive event-invariant features and thus benefit the detection of fake news on newly arrived events. It consists of three main components: a multi-modal feature extractor, the fake news detector, and the event discriminator. The multi-modal feature extractor is responsible for extracting the textual and visual features from posts. It cooperates with the fake news detector to learn the discriminating representation of news articles. The role of the event discriminator is to remove the event-specific features and keep shared features among the events.
@@
In another work (Wang et al., 2021a), Wang et al. propose MetaFEND framework which is able to detect fake news on emergent events with a few verified posts using an event adaption strategy. MetaFEND framework has two stages: event adaption and detection. In the event adaption stage, the model adapts to specific events and then in the detection stage, the event-specific parameter is leveraged to detect fake news on a given event. Although MetaFEND does not apply a generative architecture, it leverages an auxiliary network to learn event-specific parameter set to improve the efficiency of the fake news detector.
@@
The last example is a work (Silva et al., 2021b) by Silva et al where they propose a cross-domain framework using text and propagation network. The proposed model consists of two components: an unsupervised domain embedding learning; and a supervised domain-agnostic news classification. The unsupervised domain embedding exploits text and propagation network to represent a news domain with a low-dimensional vector. The classification model represents each news record as a vector using the textual content and the propagation network. Then, the model maps this representation into two different subspaces such that one preserves the domain-specific information. Later on, these two components are integrated to identify fake news while exploiting domain-specific and cross-domain knowledge in the news records.

### v7

@@ 3. Multi-modal model study
Extracted features and the way they are fused play an important role in the model architecture. In fact, model-based and feature-based studies are closely related through fusion strategies, which makes the demarcation of these two studies very difficult. Hence, in this section, we first discuss common fusion strategies as the point of connection between the two studies. Furthermore, we categorize existing works based on the machine learning techniques exploited by each work. Specifically, we classify them into two main categories: 1) classic machine learning and 2) deep learning-based solutions. In this section, we discuss each category in detail.
Early (Feature)FusionClassifierTextual FeaturesVisual FeaturesAcoustic Featuresmult-imodal FeaturesFinal Decision
Figure 2. Early fusion mechanism
Classifier 1Classifier 2Classifier nLate (Decision)FusionTextual FeaturesVisual FeaturesAcoustic FeaturesDecision 1Decision 2Decision nFinal Decision
Figure 3. Late fusion mechanism.
Early (Feature)FusionClassifierTextual FeaturesVisual FeaturesAcoustic Featuresmult-imodal FeaturesDecision AClassifier 1Classifier 2Classifier n
Late (Decision) Fusion
Late (Decision) Fusion
Textual FeaturesVisual FeaturesAcoustic FeaturesDecision 1Decision 2Decision nDecision BFinal Decision
Figure 4. A hybrid of early and late fusion mechanisms.
@@ 3.1. Fusion Mechanisms
Data fusion is the process of combining information from multiple modalities to take advantage of all different aspects of the data and extract as much information as possible to improve the performance of machine learning models, as opposed to using a single data aspect or modality. Different fusion mechanisms have been used to combine features from different modalities, including those mentioned in the previous section. Fusion mechanisms are often categorized into one of the following groups:
@@
also known as feature-level fusion, this refers to combining features from different data modalities at an early stage using an operation, which is often concatenation. This type of fusion is often performed ahead of classification. If the fusion process is done after feature extraction, it is sometimes referred to as intermediate fusion (Martínez and Yannakakis, 2014; Lahat et al., 2015; Boulahia et al., 2021).
@@
also known as decision-level or kernel-level fusion, this is usually done in the classification stage. This method depends on the results obtained by each data modality individually. In other words, the modality-wise classification results are combined using techniques such as sum, max, average, and weighted average. Most of the late fusion solutions use handcrafted rules, which are prone to human bias and are far from real-world peculiarities (Martínez and Yannakakis, 2014; Lahat et al., 2015; Boulahia et al., 2021).
3.2. Comparison of fusion mechanisms
In most cases, early fusion is a complex operation, whereas late fusion is easier to perform (Atrey et al., 2010) because, unlike early fusion where the features from different modalities (e.g., image and text) may have different representations, the decisions at the semantic level usually have the same representation. Therefore, the fusion of decisions is easier than the fusion of features. However, the late fusion strategy does not utilize the feature-level correlation among modalities, which may improve classification performance. In fact, it is shown that in many cases, the early fusion of different modalities outperforms multi-modal late fusion while applying deep learning or classic machine learning classifiers (Dong et al., 2014; Gallo et al., 2020). For instance, early fusion of images and texts while using BERT and CNN on the UPMC Food-101 dataset [^12] http://visiir.lip6.fr/explore (Wang et al., 2015) outperforms late fusion of these modalities.
Another advantage of early fusion is that it requires less computation time because training is performed only once, whereas late fusion needs multiple classifiers for local decisions (Atrey et al., 2010). However, to have the best of both worlds, there are hybrid approaches as well, which take advantage of both early and late fusion strategies (Atrey et al., 2010). Fig. 3 to Fig. illustrate simplified schemes of various fusion mechanisms for multi-modal learning4. Traditional and modern approaches for detecting multi-modal misinformation, some of which employ fusion mechanisms, are covered in the section that follows.
@@ 3.3. Classic Machine Learning Solutions
As we discussed earlier, a vast majority of misinformation detection methods leverage a single modality, a.k.a. aspect of news articles, e.g., text (Horne and Adali, 2017; Wu et al., 2017; Guacho et al., 2018; Shu et al., 2020a), image (Huh et al., 2018; Qi et al., 2019; Abdali et al., 2021a; Choudhary and Arora, 2021), user features (Wu and Liu, 2018; Shu et al., 2018, 2019b), and temporal properties (Kumar and Shah, 2018; Shu et al., 2020b; Song et al., 2021b). However, recently, there have been very few works that incorporate various aspects of a news article using classic machine learning techniques to create multi-modal article representations.
For instance, a work by Shu et al. (K. Shu and Liu, 2019) proposes individual embedded representations for text, user-user interactions, user-article interactions, and publisher-article interactions, and defines a joint optimization problem leveraging these individual representations. Finally, they apply a “Non-convex Optimization” solution via the Alternating Least Squares (ALS) algorithm to solve the proposed optimization problem.
@@
Another study (Abdali et al., 2021b) presents the K-Nearest Hyperplanes (KNH) graph, a new type of graph generalization where nodes are higher-order Euclidean subspaces formed by algebraic structures, aimed at multi-aspect modeling of news articles.
More recently, Meel et al. (Hakak et al., 2021) have proposed an “Ensemble Framework” which leverages text embedding, a score calculated by cosine similarity between image caption and news body, and noisy images. Despite the fact that some of the modules of this model, e.g., text embedding generator, leverage deep attention-based architecture, the classification process is done via a classic ensemble technique, i.e., max “Voting”.
Summarily, due to the success of deep learning-based techniques in feature extraction and classification tasks, classic machine learning-based techniques are not commonly used these days. However, considering the fact that deep learning techniques are data-hungry and require a lot of effort for training and fine-tuning the models, depending on the applications, classic machine learning techniques are still being used solely or in conjunction with deep learning techniques.
@@ 3.4. Deep Learning Solutions
Due to the impressive success of deep neural networks in feature extraction and classification of text, images, and many other modalities, they have been widely exploited by research scientists over the past few years for a variety of multi-modal tasks, including misinformation detection. We may categorize deep learning-based multi-modal misinformation detection into five categories: concatenation-based, attention-based, generative-based, graph neural network-based, and cross-modality discordance-aware architectures as demonstrated in Fig. 5. In what follows, we summarize and categorize the existing works into the aforementioned categories.
Multi-modal Misinformation Detection Classic Machine Learning Deep Learning Embeddings Concatenation Algebraic Joint Structure Ensemble Learning & Voting Non-convex Optimization Concatenation-based Networks Attention-based Networks Graph Neural Networks Cross-modal Discordance Aware Networks Foundation Models & Prompt-based Techniques
Figure 5. An overview of the multi-modal model study.
@@ 3.4.1. Concatenation-based Architectures
The majority of the existing work on multi-modal misinformation detection embeds each modality, e.g., text or image, into a vector representation and then concatenates them to generate a multi-modal representation that can be utilized for classification tasks. For instance, Singhal et al. propose using pretrained XLNet and VGG-19 models to embed text and image, respectively, and then classify the concatenation of the resulting feature vectors to detect misinformation (Singhal et al., 2019).
In another work (Segura-Bedmar and Alonso-Bartolome, 2022), Bartolome et al. exploit a Convolutional Neural Network (CNN) that takes as inputs both text and image corresponding to an article, and the outputs are concatenated into a single vector. Qi et al. extract text, Optical Character Recognition (OCR) content, news-related high-level semantics of images (e.g., celebrities and landmarks), and visual CNN features of the image. Then, in the stage of multi-modal feature fusion, text-image correlations, mutual enhancement, and entity inconsistency are merged by concatenation operation (Qi et al., 2021).
In another work (Rezayi et al., 2021), Rezayi et al. leverage network, textual, and relaying features such as hashtags and URLs and classify articles using the concatenation of the feature embeddings. Works in (Shu et al., 2019a; Raj and Meel, 2022) are other examples of this category of deep learning-based solutions.
@@ 3.4.2. Attention-based Architectures
As mentioned above, many architectures simply concatenate vector representations, thereby failing to build effective multi-modal embeddings. Such models are not efficient in many cases. For instance, the entire text of an article does not necessarily need to be false for the corresponding image and vice versa to consider the article as misinformative content. Thus, some recent works attempt to use the attention mechanism to attend to relevant parts of images, text, etc. The attention mechanism is a more effective approach for utilizing embeddings, as it produces richer multi-modal representations.
@@
For instance, a work by Sachan et al. (Sachan et al., 2021) proposes Shared Cross Attention Transformer Encoders (SCADE), which exploits CNNs and transformer-based methods to encode image and text information and utilizes cross-modal attention and shared layers for the two modalities. SCADE pays attention to the relevant parts of each modality with reference to the other.
@@
Another example is a work by Kumari et al. (Kumari and Ekbal, 2021), where a framework is developed to maximize the correlation between textual and visual information. This framework has four different sub-modules: Attention-Based Stacked Bidirectional Long Short Term Memory (ABS-BiLSTM) for textual feature representation, Attention-Based Multilevel Convolutional Neural Network–Recurrent Neural Network (ABM-CNN–RNN) for visual feature extraction, multi-modal Factorized Bilinear Pooling (MFB) for feature fusion, and finally Multi-Layer Perceptron (MLP) for classification.
@@
In another study, Qian et al. (Qian et al., 2021) introduce the Hierarchical Multi-modal Contextual Attention Network (HMCAN) architecture. This architecture leverages a pre-trained BERT and convolutional ResNet50 to create word and image embeddings. It also employs a multi-modal contextual attention network to investigate multi-modal context information. HMCAN uses various multi-modal contextual attention networks to form a hierarchical encoding network, aiming to explore and capture the rich hierarchical semantics of multi-modal data.
@@
Another example is (Jin et al., 2017), where Jin et al. fuse features from three modalities, i.e., textual, visual, and social context, using an RNN that utilizes an attention mechanism (att-RNN) for feature alignment. Jing et al. propose TRANSFAKE (Jing et al., 2021) to connect features of text and images into a series and feed them into a vision-language transformer model to learn the joint representation of multi-modal features. TRANSFAKE adopts a preprocessing method similar to BERT for concatenated text, comments, and images.
@@
In another work (Wang et al., 2022a), Wang et al. apply scaled dot-product attention on top of image and text features as a fine-grained fusion and use the fused feature to classify articles.
@@
Wang et al. propose a deep learning network for biomedical informatics that leverages visual and textual information and a semantic- and task-level attention mechanism to focus on the essential contents of a post that signal anti-vaccine messages (Wang et al., 2021b).
@@
Another example is the study by Lu et al., where they concatenate representations of user interaction, word representations, and propagation features after implementing a dual co-attention mechanism. The goal is to capture the correlations between users’ interactions/propagation and the tweet’s text (Lu and Li, 2020).
@@
Finally, Song et al. (Song et al., 2021a) propose a multi-modal fake news detection architecture based on Cross-modal Attention Residual (CARN) and Multichannel Convolutional Neural Networks (CARMN). Cross-modal Attention Residual (CARN) selectively extracts the information related to a target modality from a source modality while maintaining the unique information of the target.
@@ 3.4.3. Generative Architectures
In this category of deep learning solutions, the goal is to either apply Generative Networks or use auxiliary networks to learn individual or multi-modal representations, spaces, or parameters in order to improve the classification performance of the fake news detector.
@@
As an example, Jaiswal et al. propose a BERT-based multi-modal variational Autoencoder (VAE) (Jaiswal et al., 2021) that consists of an encoder, decoder, and a fake news detector. The encoder encodes the shared representations of both the image and text into a multidimensional latent vector. The decoder decodes the multidimensional latent vector into the original image and text, and the fake news detector is a binary classifier that takes the shared representation as an input and classifies it as either fake or real.
@@
Similarly, Kattar et al. propose a deep multi-modal variational autoencoder (MVAE) (Khattar et al., 2019) which learns a unified representation of both the modalities of a tweet’s content. Similar to the previous work, MVAE has three main components: an encoder, a decoder, and a fake news detector that utilizes the learned shared representation to predict if a news is fake or not.
@@
Like the previous work, a work by Zeng et al. (Zeng et al., 2020) proposes to capture the correlations between text and image by a VAE-based multi-modal feature fusion method. In another work, Wang et al. propose Event Adversarial Neural Networks (EANN) (Wang et al., 2018), an end-to-end framework which can derive event-invariant features and thus benefit the detection of fake news on newly arrived events. It consists of three main components: a multi-modal feature extractor, the fake news detector, and the event discriminator. The multi-modal feature extractor is responsible for extracting the textual and visual features from posts. It cooperates with the fake news detector to learn the discriminating representation of news articles. The role of the event discriminator is to remove the event-specific features and keep shared features among the events.
@@
In another work (Wang et al., 2021a), Wang et al. propose the MetaFEND framework, which is able to detect fake news on emergent events with a few verified posts using an event adaptation strategy. The MetaFEND framework has two stages: event adaptation and detection. In the event adaptation stage, the model adapts to specific events, and then in the detection stage, the event-specific parameter is leveraged to detect fake news on a given event. Although MetaFEND does not apply a generative architecture, it leverages an auxiliary network to learn an event-specific parameter set to improve the efficiency of the fake news detector.
@@
The last example is a work (Silva et al., 2021b) by Silva et al., where they propose a cross-domain framework using text and propagation network. The proposed model consists of two components: an unsupervised domain embedding learning and a supervised domain-agnostic news classification. The unsupervised domain embedding exploits text and propagation network to represent a news domain with a low-dimensional vector. The classification model represents each news record as a vector using the textual content and the propagation network. Then, the model maps this representation into two different subspaces such that one preserves the domain-specific information. Later on, these two components are integrated to identify fake news while exploiting domain-specific and cross-domain knowledge in the news records.